In last Saturday’s parade,Tom counted a total of 76 bicycles and
tricycles, for a total of 227 wheels. How many tricycles were in
the parade?

Answer by mananth(16949)

letthe number of tricycles be x
the number of bicycles will be 76-x
3x+2(76-x)=227
3x+152-2x=227
x=227-152
x=75
bicycles will be 76-75= 1